The following facts relate to Coronado Corporation.

1. Deferred tax liability, January 1, 2020, $20,200.
2. Deferred tax asset, January 1, 2020, $0.
3. Taxable income for 2020, $95,950.
4. Pretax financial income for 2020, $202,000.
5. Cumulative temporary difference at December 31, 2020, giving rise to future taxable amounts, $242,400.
6. Cumulative temporary difference at December 31, 2020, giving rise to future deductible amounts, $35,350.
7. Tax rate for all years, 20%.
8. The company is expected to operate profitably in the future.

Compute income taxes payable for 2020:

Prepare the journal entry to record income tax expense, deferred income taxes, and income taxes payable for 2020. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Prepare the income tax expense section of the income statement for 2020, beginning with the line “Income before income taxes.” (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)
Answer #1

Answer 1:

Credit Date Year 2020 | Accounts Titles and Explanation Income Tax Expense Deferred Tax Asset Income Tax Payable Deferred Tax

Working:

Income Tax Expense = Pretax financial income for 2020 * Tax rate = $202,000 * 20% = $40,4000

Income Tax Payable = Taxable income for 2020 * Tax rate = $95,950 * 20% = $19,190

Cumulative temporary difference at December 31, 2020, giving rise to future taxable amounts =$242,400

Deferred tax liability required at December 31, 2020 = 242000 * 20% = $48,480

Deferred tax liability, January 1, 2020 = $20,200

Hence: Increase in deferred tax liability in 2020 = 48480 - 20200 = $28,280

Cumulative temporary difference at December 31, 2020, giving rise to future deductible amounts, $35,350.

Deferred tax assets balance required at December 31, 2020 = 35350 * 20% = $7070

Deferred tax asset, January 1, 2020 = $0

Hence: Increase in deferred tax asset in 2020 = 7070 - 0 = $7,070

Answer 2:

Coronado Corporation Income Statement (Partial) For the Year ended December 31, 2020 $202,000 Income before income taxes Inco